August 24, 2016 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, August 24, 2016
11:00 am, EST
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at .

Calibration Tasks

  1. Spring 2016 Recon Checklist
  2. Mcsmear updates - mcsmear 20160323
  3. Data Validation
  4. GlueX Detector Performance Document

Upcoming Calibrations

For Second Production Run

  • Fix handling of individually-read-out TAGM columns (Alex B.)
  • Tagger hit merging (Nathan/Alex B.)
  • Updates to FCAL non-linearity (Adesh)
  • Recalibrate gains for early runs ~<10600 (Adesh)
  • BCAL energy non-linearity correction (Will M.)
  • FDC wire timing & reconstruction (Lubomir/Simon/Alex A.)
  • TOF ADC energy and timing calibrations [dropped] (Beni)
  • TOF alignment (Beni/Simon)
  • Beam energy calibration (Alexandre D.)
  • General timing realignment (Sean)
  • Updated error matrices
    • BCAL (Mark D.)
    • FCAL (Adesh/Matt)
    • Tracking (Simon)

For Future Production Runs

  • Add timing information to FCAL clustering/inner ring calibrations? (Adesh/Matt S.)
  • BCAL TDC calibrations (Andrei)
  • FDC per-channel calibrations (Alex A./Mike S./?)
  • TAGH alignment (Nathan)
  • Tagger energy map (Richard J.?)


  1. Announcements
    • Latest monitoring run progress
  2. Calibration Plans for Second Production Run
    • Deadline: August 31?
  3. Fall Run Calibrations
  4. Calibration Updates
  5. Simulations
    • Data/Simulation comparisons
  6. AOB


Attending: Sean (NU); Curtis, Will M., Mike S., Naomi (CMU); Simon, Alex A., Adesh, Brad, Alex B., Nathan, Mark I., Paul, Beni, Mark D., Elton (JLab)

  1. Announcements
    • The latest monitoring run has finished, and the results are up on the web. These results include the updated FDC reconstruction and new EVIO parser. There is some ongoing work to optimizing the processing and administration of the different accounts.
    • Mark Ito announced that there is a new sim-recon release.
  2. Calibration Plans for Second Production Run
    • We set a hard deadline of next Weds., August 31st, for including improvements for the next data production launch. Paul suggested prioritizing the reconstruction and analysis launch of the "golden runs" for DNP presentations, and this was generally agreed on.
    • There was a meeting last Thursday where Andrei presented his plans for precision BCAL calibrations. This was very useful, and he is currently processing the feedback from the meeting. These calibrations are useful, but not crucial so the next launch will not depend on them.
    • Per-channel FDC calibrations are also a longer term project. Alex A. and Mike S. will discuss them offline.
    • Will M. reported that he has some preliminary non-linear corrections for the BCAL as a function of energy. Corrections as a function of z are still under study. Recent studies of electron energy deposition show a different non-linear dependance than neutral showers, which need to be understood. This will be discussed in more detail in tomorrow's Calorimetry meeting, and a plan for what to include in the reconstruction launch will be determined. Note that the meeting time is shifted to 12:15 PM JLab time, due to a local seminar.
    • Mark D. plans to have updated covariance matrices based on simulation studies for this launch.
    • Simon reported that the tracking errors look generally fine, except in the transition region between the CDC and FDC, where they are small by a factor ~2.
    • For the TOF geometry, Simon reported that Beni found that the spacing between the paddles was different for the two planes. This requires some changes to the structure of the XML file. Paul offered to help with updating sim-recon to support any changes in DTOFGeometry.
    • Sean will follow up with Alexandre and Richard to see if there are any changes in the beam energy and tagger tables to include.
    • Simon is looking into some FDC pattern recognition issues which are reducing the reconstruction efficiency when there are 3 or more tracks in the forward region. It is hoped that some improvement can be made by the launch.
    • Elton asked about Kaon ID performance. Paul and Sean mentioned that they see clear signals in strangeness channels. Paul will put together a particla ID report at some future analysis meeting.
  3. Fall Run Calibrations
    • Sean is finishing up initial calibration tests with ROOT 6. Authors are in charge of verifying that their scripts still produce correct results. He has been contacting individuals about non-trivial problems that he has encountered.
    • Calibration plans were reviewed. Two stages are planned: a mostly automated stage will run offline or online, depending on the data flow and resources available. This will be decided before the run begins. The second stage will be to produce outputs (histograms and skims) for calibrations that will still be performed manually.
    • A first pass at the calibration list was performed. Sean is waiting for feedback, and he will produce a matrix for who owns what procedures/code.
    • The question of how well our calibrations could be performed depending on beam frequency and magnetic field. A solenoid current of 1350A is expected. Spring data with currents of 1200A and 1350A and beam frequencies of 250 Mhz and 500 Mhz were all well-calibrated. Sean mentioned that a constant beam frequency is desirable for ease of calibration and analysis. The beam frequency will be discussed more at the run coordination meetings, of which the first is tomorrow.
    • Naomi will contribute code for CDC gain calibrations.
    • Beni updated his plugin for the TOF calibrations. He also pointed out that in regard to the 6-fold ambiguity, his code was measuring times with respect to the time of the TOF crate, while the standard sim-recon calculation used the TS crate. For some reason, the TOF crate will sometimes shift by 2ns. This is not understood He changed his calculations to use the TS crate, which is fine except when taking cosmic data with the standalone TOF DAQ. Dealing with these two situations is a pain.
  4. Calibration Updates
    • Adesh has been working on precision FCAL timing by comparing neighboring blocks in the same shower. Using this procedure gives very bad resolutions, which is not understood. The time offsets that are obtained are very large, up ot +-20 ns, which seems unphysical. He has sent some information to Sean, and they will discuss it further offline. He is also looking into comparing shower times to RF times. Mike was looking into this before, and has some code that is close to working (just has some features that are not understood). He will communicate this to Adesh.
  5. Simulations
    • Sim 1.1 is finished. All the files are accessible from tape. Mark needs to update the webpage for this data.
    • Alex A. is running the plugins from the analysis launch over the simulated data. It is running very fast, 50% completed since yesterday, should be finished by tomorrow.
    • Will M. is in the process of comparing truth and reconstructed information in the BCAL simulation.